Jiminy's Cricket Farm just issued a 20-year, 7 percent, semiannual bond. The bond currently sells for 108 percent of its face value. What is the company’s pretax cost of debt? What is the company’s aftertax cost of debt if the tax rate is 23 percent?

Pretax cost of debt 6.29%

Aftertax cost of debt 4.84%

Please check the answers and show all work typed out. No excel or grid style please as I am on mobile.
Answer #1
Maturity Value of bonds 1000
Semi annual cash interest (1000*7%*6/12) 35
PVf at 3.145% for 40th period 0.289783
Annuity PVF at 3.145% for 40 periods 22.58242
Present value of Maturity value 289.783
Present value of Interest payments 790.3847
Issue price 1080.168
hence, Annual cost of debt = 3.145*2= 6.29%
Pre-tax cost of debt = 6.29%
After tax cost of debt = Pre-tax cost of debt (1-Tax rate)
6.29% (1-0.23) = 4.84%